Let f, g: R → R be defined by f(x) = 2x + 1 and g (x) = x^2 – 2, ∀ x ∈ R, respectively. Then, find g o f. 

Grade:12

1 Answers

Harshit Singh
2 years ago
Dear Student

Given,
f(x)=2x+1and g(x)=x^2–2,∀x∈R
Thus, g o f = g (f (x))

= g (2x + 1)
=(2x+1^)2–2
=4x^2+4x+1–2
=4x^2+4x-1

Thanks